# I Know About “Up”! Enhancing Spatial Reasoning in Visual Language Models Through 3D Reconstruction

### v1

@@ 5.2 Experimental Result
Single View vs Multiple View. Our study investigated the impact of single-view versus multi-view 3D reconstruction on the visual spatial reasoning capabilities of VLMs. We assessed two VLMs: ZeroVLM (L) and ZeroVLM (M), using both single-view and multi-view datasets. The results, shown in Figure 6, reveal that ZeroVLM excels in single-view conditions with an average accuracy of 79.28%, but its performance drops to 58.42% in multi-view conditions. ZeroVLM (M) performs less well overall, with average accuracy of 56.95% for single-view and 56.23% for multi-view. Overall, ZeroVLM (L) outperforms ZeroVLM (M) in visual spatial reasoning tasks, particularly in single-view scenarios. While ZeroVLM (M) possesses unique advantages in handling multi-modalities, there is still room for improvement in specific visual spatial reasoning tasks (Battaglia et al., 2018).
@@
L-V $\checkmark$ 87.80 87.25 88.46 75.21
R-V $\checkmark$ 86.40 88.23 90.38 83.71
Ra-V $\checkmark$ 87.00 89.21 86.53 78.33
@@
Table 3: Performance comparison of ZeroVLM (L) under different view combinations. L-V represents the left-view images, R-V represents the right-view images, Ra-V represents the random- view images, and M-V represents the multi-view images (excl. the original image). Bests are in bold, and second bests are underlined.
@@
Effect on different view prompts. The primary goal of our concluding study was to explore the potential enhancement of visual spatial reasoning abilities in VLMs by employing different view prompts. We sought to investigate if leveraging contextual connections through textual cues could significantly improve the performance of VLMs in understanding and reasoning about visual spatial relationships. To achieve this, we applied various view prompts with ZeroVLM and assessed its visual spatial reasoning abilities across both single-view and multi-view datasets. These view prompts were designed to provide contextual cues that could aid the model in interpreting visual spatial information more effectively. The results of this comprehensive assessment are presented in Table 3. The experiment demonstrated that the use of different prompts can indeed enhance the visual spatial reasoning capabilities of VLMs. Overall, our study concludes that employing a view prompt can significantly enhance the visual spatial reasoning abilities of VLMs.
@@
Case study. Our study aims to investigate the impact of 3D reconstruction on the visual spatial reasoning capabilities of VLM. We assess whether 3D reconstruction (Zero-1-to-3) enhances these capabilities from the original dataset to the single-view dataset and compare the performance of ZeroVLM (L) and ZeroVLM (M) using single-view images after 3D reconstruction. Comparative analysis focuses on the visual spatial reasoning capabilities of ZeroVLM on both datasets. Figure 7 illustrates the enhancement by ZeroVLM using the single-view dataset. Figure 8 presents the performance comparison between ZeroVLM (L) and ZeroVLM (M). Results show that ZeroVLM (L) has improved visual spatial reasoning with the single-view dataset after 3D reconstruction and outperforms ZeroVLM (M). This suggests that 3D reconstruction provides useful spatial information aiding visual reasoning.
@@
Figure 7: Different results were obtained by ZeroVLM when evaluating datasets with and without 3D reconstruction. View prompts are omitted to save space.
@@
Figure 8: Different results were obtained when inputting the 3D reconstructed datasets into ZeroVLM (L) and ZeroVLM (M). View prompts are omitted to save space.
6 Conclusion
@@
This study presents a novel model, called ZeroVLM, for enhancing the visual spatial reasoning capability of VLMs through 3D reconstruction. In particular, ZeroVLM use the Zero-1-to-3 to generate different views over different angles of the input image, and generate answers with a specially designed view prompt. We validate the effectiveness of our VLMs by comparative experiments from various views, including single and multiple viewpoints, and evaluated the performance of LLaVA and MiniGPT-4 under our ZeroVLM. The experimental results suggest that 3D reconstruction with a view prompt from a single perspective can effectively optimize the model’s visual spatial reasoning ability. Future research could also develop models that dynamically adjust viewpoints based on task requirements and integrate additional modal information, such as video and audio, to enhance multi-modal processing capability.
7 Limitations
@@ Appendix A Appendix
Figure 9: The primary difference between LLaVA and ZeroVLM (L) lies in their input handling. LLaVA processes a single image as input, whereas ZeroVLM (L) is designed to handle multiple images. Due to the limited expressive capacity of individual images, we restrict the input for ZeroVLM (L) to two images.
Figure 10: The 3D reconstructed single-view and multi-view versions of each dataset are presented.

### v2

@@ 5.2. Experimental Result
Single View vs Multiple View. In this experiment, we investigated the impact of single-view versus multi-view 3D reconstruction on the visual spatial reasoning capabilities of VLMs. We assessed two VLMs: ZeroVLM (L) and ZeroVLM (M), using both single-view and multi-view datasets. The results, shown in Figure 6, reveal that ZeroVLM excels in single-view conditions with an average accuracy of 79.28%, but its performance drops to 58.42% in multi-view conditions. ZeroVLM (M) performs less well overall, with average accuracy of 56.95% for single-view and 56.23% for multi-view. Overall, ZeroVLM (L) outperforms ZeroVLM (M) in visual spatial reasoning tasks, particularly in single-view scenarios. While ZeroVLM (M) possesses unique advantages in handling multi-modalities, there is still room for improvement in specific visual spatial reasoning tasks (Battaglia et al., 2018). The reason that multi-view models failed might be due to the fact that there are too many elements in the multi-view image dataset and the model cannot distinguish them correctly.
@@
L-V $\checkmark$ 89.20 88.63 88.62 76.32
R-V $\checkmark$ 87.60 89.41 90.38 84.21
Ra-V $\checkmark$ 88.40 90.09 87.42 79.56
@@
Table 3. Performance comparison of ZeroVLM (L) under different view combinations. L-V represents the left-view images, R-V represents the right-view images, Ra-V represents the random- view images, and M-V represents the multi-view images (excl. the original image). Bests are in bold, and second bests are underlined.
@@
Effect on different view prompts. In this experiment, we aim to explore the potential enhancement of visual spatial reasoning abilities in VLMs by employing different view prompts. We sought to investigate if leveraging contextual connections through textual cues could significantly improve the performance of VLMs in understanding and reasoning about visual spatial relationships. To achieve this, we applied various view prompts with ZeroVLM and assessed its visual spatial reasoning abilities across both single-view and multi-view datasets. These view prompts were designed to provide contextual cues that could aid the model in interpreting visual spatial information more effectively. The results of this comprehensive assessment are presented in Table 3. The experiment demonstrated that the use of different prompts can indeed enhance the visual spatial reasoning capabilities of VLMs. Overall, our study concludes that employing a view prompt can enhance the visual spatial reasoning abilities of VLMs.
@@
Case study. Our study aims to investigate the impact of 3D reconstruction on the visual spatial reasoning capabilities of VLM. We assess whether 3D reconstruction (Zero-1-to-3) enhances these capabilities from the original dataset to the single-view dataset and compare the performance of ZeroVLM (L) and ZeroVLM (M) using single-view images after 3D reconstruction. The comparative analysis focuses on the visual spatial reasoning capabilities of ZeroVLM on both datasets. Figure 7 illustrates the enhancement by ZeroVLM using the single-view dataset. Figure 8 presents the performance comparison between ZeroVLM (L) and ZeroVLM (M). Results show that ZeroVLM (L) has improved visual spatial reasoning with the single-view dataset after 3D reconstruction and outperforms ZeroVLM (M). This suggests that 3D reconstruction provides useful spatial information aiding visual reasoning.
@@
Figure 7. Different results were obtained by ZeroVLM when evaluating datasets with and without 3D reconstruction. View prompts are omitted to save space.
@@
Figure 8. Different results were obtained when inputting the 3D reconstructed datasets into ZeroVLM (L) and ZeroVLM (M). View prompts are omitted to save space.
6. Conclusion
@@
This work studies the task of visual spatial reasoning and presents a simple and novel model, called ZeroVLM, for enhancing the visual spatial reasoning capability of VLMs through 3D reconstruction. In particular, ZeroVLM uses the Zero-1-to-3 model to generate different views over different angles of the input image, and uses VLMs to generate spatial reasoning answers with the stitched view image and a specially designed view prompt. We validate the effectiveness of our VLMs by comparative experiments from various views, including single-view and multiple-view images, and evaluated the performance of LLaVA and MiniGPT-4 under our ZeroVLM. The experimental results suggest that 3D reconstruction with a view prompt from a single perspective can effectively enhance the model’s visual spatial reasoning ability. Future research could also develop models that dynamically adjust view prompts based on task requirements and integrate additional modal information, such as video and audio, to enhance multi-modal processing capability.
7. Limitations
@@ Appendix A Appendix
Figure 9. The primary difference between LLaVA and ZeroVLM (L) lies in their input handling. LLaVA processes a single image as input, whereas ZeroVLM (L) is designed to handle multiple images. Due to the limited expressive capacity of individual images, we restrict the input for ZeroVLM (L) to two images.
Figure 10. The 3D reconstructed single-view and multi-view versions of each dataset are presented.